# Learning Noise-Robust Stable Koopman Operator for Control with Physics-Informed Observables

### v1

@@ I Introduction
Data-driven methods have shown promising potential in modeling and control of nonlinear systems [1, 2]. Interested readers are encouraged to read several excellent reviews of Koopman operator [1, 3, 4, 5, 6]. The Koopman operator has been applied across diverse fields, including robust COVID-19 prediction [7], reinforcement learning [8], soft robotics control [9], stability analysis of genetic toggle switches [10], and modeling turbulent shear flows [11].
@@
The biggest benefit of Koopman operator for dynamics and control community is to facilitate the use of well-established linear control/observer techniques for the nonlinear systems[12, 13, 14], provided that a good approximation of global linear representation from data is learned. For example, model predictive control (MPC) is typically computationally expensive due to the need to solve non-convex optimization online. Traditional linearization techniques that approximate the nonlinear system around a local reference point might result in suboptimal MPC performance and could struggle with significant deviations from this point. On the other hand, the linear representation from Koopman operator can be used to transform the non-convex optimization of the nonlinear MPC into convex optimization of a linear MPC as shown by [15]. In addition to classical MPC, such linear representation can also help LQR [13], anomaly detection [16] and fault tolerant control [17] to nonlinear systems.
Despite the several successful application of Koopman operator, there still remain several challenges. Colbrook et al. [18] examine the limitations of the Koopman operator and systematically determines when and how Koopman spectral properties can be learned. Wu et al. [19] reveal that under mild noise and nonlinearity, DMD can struggle to accurately recover and predict dynamics, even with an exact linear model.
@@
Among various challenges, the first challenge is finding an appropriate invariant subspace for the Koopman operator. This involves selecting suitable basis functions that span an invariant subspace. Common choices of basis functions include radial basis functions (RBFs) [20] or polynomial functions [21]. Another approach is to learn the observables from dictionary of functions [22], or identifying dominant eigenfunctions from data [13]. Otto et al. [23] formulate the dynamics of observables under the Koopman generator as a bilinear hidden Markov model. They determine the model parameters using the expectation-maximization (EM) algorithm, thereby eliminating the need to explicitly specify a dictionary of functions, as required by existing methods like Extended Dynamic Mode Decomposition (EDMD). However, in order to lift the system, one has to perform a non-trivial joint optimization of hidden state and dynamics.
Recently, the concept of Polyflow was introduced by [24]. Polyflow employs the governing equations of the system to construct basis functions, offering a non-local linearization approach for nonlinear differential equations. This concept, which we refer to as physics-informed observables, is closely related to the derivative-based approximation of Koopman operator [25] except closure is not ignored in Polyflow. Despite the success of initial attempt of Polyflow for MPC [26], simply minimizing the one-step ahead error will be prone to noisy data.
The second challenge in learning Koopman operator from data is ensuring long-term prediction accuracy. The use of roll-out techniques for recurrent loss has proven beneficial when data is limited and noisy both in the deep learning community [27, 28, 29] and the reduced-order modeling community [30]. However, optimizing the recurrent loss can be tricky and prone to poor local minimum in practice.
The third challenge in Koopman operator-based control is ensuring the stability of the Koopman operator itself since the real part of any continuous time eigenvalue of the Koopman operator of any physical system should not be positive [29, 31]. When the forward pass becomes unstable, the backward pass will be unstable as well. Lastly, curating training data for Koopman operator-based control often depends on ad-hoc selections of input signals and initial conditions, which can result in insufficient sampling of the high-dimensional phase space and, consequently, poor control performance.
@@ II-A Lifted Linear Predictor
where $f(\cdot)$ represents a generally nonlinear transition mapping, and $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ denote the system’s state and input, respectively. The objective is to predict the trajectory of the system, given an initial condition $x_{0}$ at time $t_{0}$, and a sequence of inputs $\{u_{i}\}_{i=0}^{N_{p}}$, where $N_{p}$ is the prediction horizon length. To achieve this, we define the lifted linear predictor as a dynamical system:
@@
The predictor can also take the form of a bilinear system:
$z_{k+1}$ $=Az_{k}+(Bz_{k})u_{k},$ (4)
@@
which can capture more complex dynamical behaviors. While our framework can be extended to accommodate the bilinear form, we avoid this for simplicity and to maintain readiness for application in linear control systems.
@@ II-B Koopman Operator for Controlled System
To provide a rigorous justification for constructing a linear predictor via state-space lifting, we briefly outline the Koopman operator approach for analyzing controlled dynamical systems. We will skip the discussion for Koopman Operator for autonomous system.
We consider the following discrete-time nonlinear system with control:
@@
where, $x\in\mathbb{R}^{n}$ is the state of the system and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input. $\mathcal{U}$ is the space of the available control inputs. $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the transition mapping. $x^{+}$ is the system state of the next time step.
We define the extended state comprising of the state $x$ of system and the control sequence $\mathbf{u}\in\ell(\mathcal{U})$, where $\ell(\mathcal{U})$ is the space of all sequences $(u_{i})_{i=0}^{\infty}$ with $(u_{i})_{i=0}^{\infty}\in\mathcal{U}$,
@@
where $\mathcal{S}$ is the left shift operator, i.e., $(\mathcal{S}\mathbf{u})(i)=u(i+1)$, and $u(i)$ denotes the $i$th element of the sequence $\mathbf{u}$.
The Koopman operator $\mathcal{K}:\mathcal{H}\to\mathcal{H}$ associated to (6) is defined by
@@
for any observable $\psi:\mathbb{R}^{n}\times\ell(\mathcal{U})\to\mathbb{R}$ belonging to some space of observables $\mathcal{H}$.
Despite Koopman operator $\mathcal{K}$ being a linear operator fully describing the time evolution of any observable function on the non-linear dynamical system, it is inherently infinite dimensional that makes learning Koopman operator challenging.
@@ II-C Extended DMD for Control with Physics-Informed Observables
For practical purposes, we adapt to the Extended Dynamic Mode Decomposition (EDMD) [32] approach followed by Korda and Mezić [15]. The EDMD provides a straightforward method to approximate the Koopman operator from data, thereby obtaining the lifted linear predictor in (2). Since our goal is not to predict the control sequence, the observable function will depend only on the system states. Thus, the process begins by selecting $N_{\phi}$ observables $\phi_{1},\ldots,\phi_{N_{\phi}}$ and forming the predictor’s state $z$ from the measured state $x$ of the system (1) as follows:
@@
To select suitable basis functions $\Phi$ that span an invariant subspace, we are motivated by Polyflow basis functions defined in [26]. When the dynamics $f(x,u)$ is known, we can use a special set of basis functions $\Phi(x)=T_{N}(x)$, known as Polyflow basis functions of order $N\in\mathbb{Z}^{+}$ which is defined,
@@
Using these polyflow basis functions to lift the system state, we can derive the following system, which aligns with equation (2)
@@
Time delay embedding [33] and Polyflow offer distinct approaches to state lifting. Time delay embedding relies on observing and recording the physical system’s state over time to embed the state vector. In contrast, Polyflow directly utilizes the governing equations of the system, enabling an instantaneous embedding of the state without accessing the physical system itself. Our framework assumes the existence of a governing equation, but for systems with unknown dynamics, methods such as SINDy [34] or Neural ODE [35] can be employed to develop a working governing equation.
@@ II-D Recurrent Roll-out Loss
To enhance long-term prediction accuracy, we implement a recurrent roll-out loss instead of using one-step prediction loss (8). For this, we consider a set of $M$ trajectories, each consisting of $R$ steps. Let $x_{i,j}$ denote the state at the $i$-th step of the $j$-th trajectory, where $i=0,1,\ldots,R$ and $j=1,2,\ldots,M$. The following optimization problem is then formulated:
@@
where, $A$ and $B$ are the matrices to be determined, and $u_{i}^{j}$ denotes the input at the $i$-th timestep of the $j$-th trajectory.
To determine an appropriate basis size $N$, we solve the optimization problem (13) iteratively, increasing $N$ until the approximation error is sufficiently small.
@@ II-E Curriculum Learning
Curriculum learning is a training strategy where a model is first trained on simpler data points and then gradually introduced to more challenging ones. This approach helps the model to progressively learn and adapt to increasing complexity.
@@ II-F Stabilized Formulation
To promote stability of the lifted system, we enforce stable eigenvalues of $A$ by employing the fact that a skew symmetric matrix with negative real valued diagonal elements have stable or neutral stable eigenvalues in continuous time dynamical system. For any $Q\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ matrix and $C\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ diagonal matrix with negative real diagonal elements, $\tilde{A}=Q-Q^{T}+C$ has eigenvalues in the negative real part of the complex plane. We compute the discrete-time $\hat{A}\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ by taking the matrix exponentiation of $\tilde{A}=Q-Q^{T}+C$ multiplied with the sampling timestep $\Delta t$.
@@
A similarity transform of $\hat{A}$ via the coordinate transform matrix $P\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ is considered to find the Koopman Operator $A$.
@@
This formulation ensures the stability of the Koopman Operator throughout the training. We name this Dissipative Koopman Operator. In contrast, a Standard Koopman Operator Matrix can be any matrix $A$ without any stability constraint.
@@ II-G Model Predictive Control
The application of lifted linear predictor within MPC is referred to as Koopman MPC (KMPC) in [15]. Here, we present the main formulation and omit the details, which can be found in the referenced literature. At each discrete time step, the MPC solves the following optimization problem over the prediction horizon $N_{p}$:
@@ II-H Iterative Learning
Models trained on trajectories that are generated by randomly sampling the phase space and control input can perform poorly when subjected to the optimal feedback control inputs. This is primarily because these trained models lack exposure to the desired scenarios during the training process, especially for relatively high-dimensional system. To address this issue, we propose an iterative data augmentation technique based on collecting response of the true system under control input that comes from the data-driven model. Such response together with control input is used to augment the original dataset to further refine and improve the model. Our rationale is that such experience from the true system can be used to better explore the phase space of nonlinear system than merely ad-hoc random sampling the phase space, which is subject to the curse of dimensionality.
Our proposed framework is named as STIRK (Stable Training with Iterative Roll-out Koopman Operator). A summary of the proposed framework is illustrated in the figure 1.
@@ III Numerical Experiments
In this section, we will evaluate our proposed framework by considering the following two tasks:
@@ III-A Prediction Performance
In order to evaluate the prediction performance of the proposed framework, we compare it with the following versions of DMD:
@@
To evaluate the prediction performance of the different algorithms, we use a normalized error metric that quantifies the discrepancy between the predicted trajectories and the ground truth trajectories. The evaluation is conducted as follows:
Let $\mathbf{x}(t)\in\mathbb{R}^{n}$ represent the true state of the system at time $t$ and $\hat{\mathbf{x}}(t)\in\mathbb{R}^{n}$ represent the predicted state at time $t$. The normalized prediction error over the trajectory is computed using the following formula:
@@
Figure 3(a) compares the reconstruction performance of various Dynamic Mode Decomposition (DMD)-based methods when trained on a single trajectory of the Van Der Pol system at a noise level of 0.0599. The legend lists the normalized error for each method. The Dissipative Koopman model with Polyflow basis functions (DISSIPATIVE-Polyflow) achieved the lowest error of $2.98\times 10^{-2}$, indicating superior accuracy. OptDMD, using both identity and Polyflow basis functions, performed similarly well, closely approaching the performance of the Dissipative Koopman model. However, since OptDMD also optimizes the initial condition, it shows a deviation from the actual initial condition. Other methods exhibited higher errors, indicating less accurate reconstructions under the given noise conditions. Figure 3(b) presents the mean squared error (MSE) trajectory over 100 timesteps for the Van Der Pol system at a noise level of 0.0599 for 10 different training sets. Each line represents the mean of the errors for a different DMD-based method, with shaded areas indicating the standard deviation of the error. The Dissipative Koopman model with Polyflow basis functions consistently maintains the lowest MSE across all timesteps, highlighting its robustness and accuracy under noisy conditions. The eDMD with Polyflow basis functions and the eDMD with RBF basis functions exhibit higher MSEs, particularly in the early timesteps, indicating less accurate predictions compared to the Dissipative model. The standard DMD with identity observables (DMD-Identity) also shows higher error compared to the Dissipative Koopman model but performs better than some of the other methods. The figures demonstrate that the Dissipative Koopman model with Polyflow basis functions outperforms other DMD-based methods in terms of prediction accuracy for the Van Der Pol system under noisy conditions. This highlights the effectiveness of incorporating dissipative properties and structured observables in enhancing the robustness and accuracy of the model.
(a) Single Trajectory Prediction for Van Der Pol System with Noise Level 0.0599. The plot compares the ground truth trajectory (black solid line) with the predictions from various DMD-based methods.
(b) Mean Squared Error (MSE) Trajectory over Timesteps for Van Der Pol System with Noise Level 0.0599. The plot shows the evolution of MSE over 100 timesteps for different DMD-based methods. Shaded areas represent mean $\pm$ standard deviation of the mean squared errors. The Dissipative Koopman model with Polyflow basis functions maintains the lowest error throughout the timesteps, demonstrating robust performance under noisy conditions.
Figure 3: Single Trajectory Prediction Performance For Van Der Pol System
@@
Figure 4: Multiple Trajectory Error Comparison for VDP system

### v2

@@ Abstract
Index Terms:
Machine learning, Koopman operator, optimal control, model predictive control
@@ I Introduction
Data-driven methods have shown promising potential in modeling and control of nonlinear systems [1, 2]. The Koopman operator, as initially proposed by Bernard Koopman [3], plays a pivotal role by providing a global linear representation of nonlinear dynamical systems [4, 5]. Interested readers are encouraged to read several excellent reviews of Koopman operator [1, 6, 7, 8, 9]. The Koopman operator has been applied across diverse fields, including robust COVID-19 prediction [10], reinforcement learning [11], soft robotics control [12], stability analysis of genetic toggle switches [13], and modeling turbulent shear flows [14].
@@
The biggest benefit of Koopman operator for dynamics and control community is to facilitate the use of well-established linear control/observer techniques for the nonlinear systems [15, 16, 17], provided that a good approximation of global linear representation from data is learned. For example, model predictive control (MPC) is typically computationally expensive due to the need to solve non-convex optimization online. Traditional linearization techniques that approximate the nonlinear system around a local reference point might result in suboptimal MPC performance and could struggle with significant deviations from this point. On the other hand, the linear representation from Koopman operator can be used to transform the non-convex optimization of the nonlinear MPC into convex optimization of a linear MPC as shown by [18, 19]. In addition to MPC, such linear representation from Koopman operator can also help LQR [16], anomaly detection [20] and fault tolerant control [21] to nonlinear systems.
@@
Despite the several successful application of Koopman operator, there still remain several challenges. The first challenge is to find an appropriate invariant subspace for the Koopman operator. This involves selecting suitable basis functions that span an invariant subspace. Common choices of basis functions include radial basis functions (RBFs) [22] or polynomial functions [23]. Interestingly, Otto et al. [24] chose to learn the latent linear model and hidden state jointly, thereby eliminating the need to explicitly specify a dictionary of functions, as required by existing methods like Extended Dynamic Mode Decomposition (EDMD). However, one has to perform a non-trivial optimization for the linear dynamics and hidden states, which are unique for each single trajectory.
Recently, the concept of Polyflow was introduced by [25]. Polyflow employs the governing equations of the system to construct basis functions, offering a non-local linearization approach for nonlinear differential equations. This concept, which we refer to as physics-informed observables, is closely related to the derivative-based approximation of Koopman operator [26] except closure is considered. Despite the success of initial attempt of Polyflow for MPC [27], simply minimizing the one-step ahead error will nevertheless be prone to noisy data. The second challenge is to ensure long-term prediction accuracy. Wu et al. [28] reveal that under mild noise and nonlinearity, DMD can struggle to accurately predict the dynamics. Although the use of roll-out techniques for recurrent loss has proven beneficial when data is limited and noisy both in the deep learning [29, 30, 31] and the reduced-order modeling community [32], optimizing the recurrent loss can be tricky and prone to poor local minimum in practice. The third challenge in Koopman operator-based control is to ensure the stability of the Koopman operator itself since the real part of any continuous-time eigenvalue of the Koopman operator of any physical system should not be positive [31, 33]. If the forward pass becomes unstable, the backward pass will be unstable as well. Lastly, curating training data for Koopman operator-based control often depends on ad-hoc selections of input signals and initial conditions, which can result in insufficient sampling especially in the high-dimensional phase space and, consequently, poor control performance. Besides, it is important to note that we do not attempt to address other challenges such as the convergence of Koopman operator. Interested readers are referred to this excellent paper by Colbrook and Mezić [34].
@@ II-A Lifted Linear Predictor
where $f(\cdot)$ represents a generally nonlinear mapping, and $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ denote the state and input of the system, respectively. The objective is to predict the trajectory of the system, given an initial condition $x_{0}$ at time $t_{0}$, and a sequence of inputs $\{u_{i}\}_{i=0}^{N_{p}}$, where $N_{p}$ is the prediction horizon length of interest. To achieve this, we define the lifted linear predictor as a linear dynamical system determined by a triplet of matrices $(A,B,C)$:
@@
It is important to note that under mild assumpesions, the predictor could also take a bilinear form:
$z_{k+1}$ $=Az_{k}+Bz_{k}u_{k},$ (4)
@@
as noted in several works [35, 36, 37, 38]. While it is natural to extend our work to accommodate the bilinear form, we avoid this for simplicity and maintain readiness for application in linear control systems.
@@ II-B Koopman Operator for Controlled System
To provide a rigorous justification for constructing a linear predictor via state-space lifting, we briefly outline the Koopman operator approach for analyzing controlled dynamical systems. We consider the following discrete-time nonlinear system with control:
@@
where, $x\in\mathbb{R}^{n}$ is the state of the system and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input. $\mathcal{U}$ is the space of the available control inputs. $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the transition mapping. $x^{+}$ is the system state of the next time step. We define the extended state comprising of the state $x$ of system and the control sequence $\mathbf{u}\in\ell(\mathcal{U})$, where $\ell(\mathcal{U})$ is the space of all sequences $(u_{i})_{i=0}^{\infty}$ with $(u_{i})_{i=0}^{\infty}\in\mathcal{U}$,
@@
where $\mathcal{S}$ is the left shift operator, i.e., $(\mathcal{S}\mathbf{u})(i)=u(i+1)$, and $u(i)$ denotes the $i$th element of the sequence $\mathbf{u}$. The Koopman operator $\mathcal{K}:\mathcal{H}\to\mathcal{H}$ associated to eq. 6 is defined by
@@
for any observable $\psi:\mathbb{R}^{n}\times\ell(\mathcal{U})\to\mathbb{R}$ belonging to some space of observables $\mathcal{H}$. Despite Koopman operator $\mathcal{K}$ being a linear operator fully describing the time evolution of any observable function on the non-linear dynamical system, it is inherently infinite dimensional that makes learning Koopman operator challenging.
@@ II-C Extended DMD for Control with Physics-Informed Observables
For practical purposes, we adapt to the Extended Dynamic Mode Decomposition (EDMD) [39] approach followed by Korda and Mezić [18] for control. The EDMD provides a straightforward method to approximate the Koopman operator from data, thereby obtaining the lifted linear predictor in eq. 2. Since our goal is not to predict the control sequence, the observable function will depend only on the system states. Thus, the process begins by selecting $N_{\phi}$ observables $\phi_{1},\ldots,\phi_{N_{\phi}}$ and forming the lifted state $z$ from the measured state $x$ of the system (1) as follows:
@@
To select suitable basis functions $\Phi$ that span an invariant subspace, we are motivated by Polyflow basis functions from Wang and Jungers [27]. When the dynamics $f(x,u)$ is known, we can use a special set of basis functions $\Phi(x)=T_{N}(x)$, known as Polyflow basis functions of order $N\in\mathbb{Z}^{+}$ which is defined,
@@
Using these Polyflow basis functions to lift the system state, we can derive the following system, which aligns with equation eq. 2
@@
Despite the similarity, time delay embedding [40] and Polyflow offer slightly distinct approaches to state lifting. Time delay embedding relies on observing and recording the physical system’s state over time to embed the state vector. In contrast, Polyflow directly utilizes the governing equations of the system, enabling an instantaneous embedding of the state without accessing the history of the physical system. Our framework assumes a known governing equation, but for systems with unknown dynamics, methods such as SINDy [41] or Neural ODE [42] can be employed as a surrogate of the governing equation.
@@ II-D Recurrent Roll-out Loss
To enhance long-term prediction accuracy, we implement a recurrent roll-out loss instead of using one-step prediction loss eq. 8. For this, we consider a set of $M$ trajectories, each consisting of $R$ steps. Let $x_{i,j}$ denote the state at the $i$-th step of the $j$-th trajectory, where $i=0,1,\ldots,R$ and $j=1,2,\ldots,M$. The following optimization problem is then formulated:
@@
where, $A$ and $B$ are the matrices to be determined, and $u_{i}^{j}$ denotes the input at the $i$-th timestep of the $j$-th trajectory. To determine an appropriate basis size $N$, we solve the optimization problem (13) iteratively, increasing $N$ until the approximation error is sufficiently small.
@@ II-E Curriculum Learning
Curriculum learning [43] is a training strategy where a model is first trained on simpler data points and then gradually introduced to more challenging ones. This approach helps the model to progressively learn and adapt to increasing complexity.
@@ II-F Stabilized Formulation
To promote stability of the lifted system, we enforce stable eigenvalues of $A$ by employing the fact that a skew symmetric matrix with negative real valued diagonal elements have stable or neutral stable eigenvalues in continuous time dynamical system [31]. For any $Q\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ matrix and $C\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ diagonal matrix with negative real diagonal elements, $\tilde{A}=Q-Q^{T}+C$ has eigenvalues in the negative real part of the complex plane. We compute the discrete-time $\hat{A}\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ by taking the matrix exponentiation of $\tilde{A}=Q-Q^{T}+C$ multiplied with the sampling timestep $\Delta t$.
@@
Note that a similarity transform of $\hat{A}$ via the coordinate transform matrix $P\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ is required to recover $A$.
@@
This formulation ensures the stability of the Koopman Operator throughout the training, which we refer as Dissipative Koopman Operator. In contrast, we refer Standard Koopman Operator as lifted system with an unconstrained matrix $A$.
@@ II-G Model Predictive Control
One of the most important applications of lifted linear predictor is to provide a surrogate model for MPC, which is referred as Koopman-MPC (KMPC) [18]. At each discrete time step, the MPC solves the following optimization problem over the prediction horizon $N_{p}$:
@@ II-H Iterative Learning
Models trained on trajectories that are generated by randomly sampling the phase space and control input could perform poorly when subjected to the optimal feedback control inputs. This is primarily because these trained models lack exposure to the desired scenarios during the training process, especially for relatively high-dimensional system. To address this issue, we propose a straightforward iterative data augmentation technique inspired by the work of Uchida and Duraisamy [44]. The idea is to iteratively collect response of the true system under control input synthesized with the previous data-driven model. Such response together with control input is used to augment the original dataset to further refine and improve the data-driven model. Our rationale is that such experience from the true system can be used to better explore the phase space of nonlinear system than merely ad-hoc random sampling the phase space, which is subject to the curse of dimensionality. Hence, our proposed framework is named as STIRK (Stable Training with Iterative Roll-out Koopman Operator). A summary of the proposed framework is illustrated in the figure 1.
@@ III Numerical Experiments
In this section, we will evaluate our proposed framework against other popular frameworks by considering the following two tasks:
@@ III-A Prediction Performance
We compare the proposed framework with the following versions of DMD:
@@
To evaluate the prediction performance of the different algorithms, we use a normalized error metric that quantifies the discrepancy between the predicted trajectories and the ground truth trajectories. Let $\mathbf{x}(t)\in\mathbb{R}^{n}$ represent the true state of the system at time $t$ and $\hat{\mathbf{x}}(t)\in\mathbb{R}^{n}$ represent the predicted state at time $t$. The normalized prediction error over the trajectory is computed using the following formula:
@@
Figure 3a compares the reconstruction performance of various Dynamic Mode Decomposition (DMD)-based methods when trained on a single trajectory of the Van der Pol system at a noise level of 0.0599. The legend lists the normalized error for each method. The Dissipative Koopman model with Polyflow basis functions (DISSIPATIVE-Polyflow) achieved the lowest error of $2.98\times 10^{-2}$, indicating superior accuracy. OptDMD, using both identity and Polyflow basis functions, performed similarly well, closely approaching the performance of the Dissipative Koopman model. However, since OptDMD also optimizes the initial condition, it shows a deviation from the actual initial condition. Other methods exhibited higher errors, indicating less accurate reconstructions under the given noise conditions. Figure 3b presents the mean squared error (MSE) trajectory over 100 timesteps for the Van der Pol system at a noise level of 0.0599 for 10 different training sets. Each line represents the mean of the errors for a different DMD-based method, with shaded areas indicating the standard deviation of the error. The Dissipative Koopman model with Polyflow basis functions consistently maintains the lowest MSE across all timesteps, highlighting its robustness and accuracy under noisy conditions. The eDMD with Polyflow basis functions and the eDMD with RBF basis functions exhibit higher MSEs, particularly in the early timesteps, indicating less accurate predictions compared to the Dissipative model. The standard DMD with identity observables (DMD-Identity) also shows higher error compared to the Dissipative Koopman model but performs better than some of the other methods. The figures demonstrate that the Dissipative Koopman model with Polyflow basis functions outperforms other DMD-based methods in terms of prediction accuracy for the Van der Pol system under noisy conditions. This highlights the effectiveness of incorporating dissipative properties and structured observables in enhancing the robustness and accuracy of the model.
Figure 3: Comparison of model prediction performance for Van der Pol system. (a) Single Trajectory Prediction for Van der Pol System with Noise Level 0.0599. (b) Mean Squared Error (MSE) Trajectory over Timesteps for Van der Pol System with Noise Level 0.0599. The plot shows the evolution of MSE over 100 timesteps for different DMD-based methods. Shaded areas represent mean $\pm$ standard deviation of the mean squared errors. The Dissipative Koopman model with Polyflow basis functions maintains the lowest error throughout the timesteps, demonstrating robust performance under noisy conditions.
@@
Figure 4: Multiple Trajectory Error Comparison for the Van der Pol (VDP) system. (a) Training Error. (b) Test Error.